Suppose the following table given below describes the production possibilities of RICE and THREADS for FINLAND and JAPAN:

 

RICE (Kg's/hour) THREADS (Meters/hour)

FINLAND 15 20

JAPAN 20 70


  1. Without trade, what is the relative price of RICE in terms of cloth in FINLAND?

What is the price in JAPAN?

  1. Which country has an absolute advantage in the production of each good?
  2. Which country has a comparative advantage in the production of each good?
  3. If both countries trade with each other, which good will each export?
  4. What is the range of prices at which trade can occur?
1
Expert's answer
2021-04-01T19:03:56-0400

1.

"Pr=\frac{Px}{Py}"

Pr= relative price

Px= price of the first product

Py= price of the second product

relative price of rice in Finland

"Pr=\frac{15}{20}=0.75"

relative price of rice in japan

"Pr=\frac{20}{17}=1.176"

2.

Finland has absolute advantage in production of threads.

Japan has absolute advantage in production of of rice.

3.

opportunity cost of one unity of rice in

Finland"=\frac{20}{15}=1.3"

1kg/hour of rice=0.85 threads in meters/hour

Japan"=\frac{17}{20}=0.85"

1kg/hour of rice=o.85 threads in meters/hour

opportunity cost of one unit of threads in

Finland"=\frac{15}{20}=0.5"

1 meter /hour of thread=0.5 kg/hour of rice

Japan"=\frac{20}{17}=1.176"

1 meter/hour of thread =1.176 kg/hour of rice

Finland has comparative advantage in production of thread.

japan has a comparative advantage of production of in rice.

4.

Finland will export threads.

Japan will export rice.

5.

trade can occur at any price between 1 and 2 kg/hour of rice per meter/hour of thread.